A circular dartboard has a radius of 20cm. Its bull's eye has a radius of 4cm. If an amateur throws a dart and it hits the board, what is the probability that the dart hits the bulls eye? answer as a %

Answer by checkley71(8403)

AREA(A)=3.14(20/2)^2
A=3.14(10)^2
A=3.14*100
A=314 AREA OF THE WHOLE DART BOARD.
B=3.14(4/2)^2
B=3.14(2^2)
B=3.14*4
B=12.56 ANSWER FOR THE BULL'S EYE.
12.56/314=.04 OR 4% CHANCE FOR A BULL'S EYE.